on a single draw from a shuffled standard deck of 52 cards, find the probability of obtaining a red face card and a spade. Please show me your work on this.
Answer by stanbon(57979)

on a single draw from a shuffled standard deck of 52 cards, find the probability of obtaining a red face card and a spade.
-----
P(red face card) = 6/52 = 3/26
------------
P(spade) = 13/52 = 1/4
==========================
Cheers,
Stan H.